OpenStudy (blockcolder):

Evaluate: \[\Large \frac{(10^4+324)(22^4+324)(34^4+324)(46^4+324)(58^4+324)}{(4^4+324)(16^4+324)(28^4+324)(40^4+324)(52^4+324)}\] So I cheated and plugged this in the calculator which gave 373 but I want to know how to evaluate this without a calculator.

OpenStudy (lgbasallote):

uhmm you just...

OpenStudy (blockcolder):

I thought you had a solution. :|

OpenStudy (turingtest):

did you try to write things as multiples of 2 ? (just brainstorming...)

OpenStudy (lgbasallote):

seems like a case of \[\large \frac{[n^4 + 324][(n+12)^4 + 324][(n+24)^4 + 324]...[(n+48)^4 + 324]}{[m^4 +324)][(m+12)^4 + 324][(m+24)^24 + 324]...[(m+48)^4 + 324]}\]

OpenStudy (turingtest):

but doesn't it seems suspicious that all the bases are even? I wanna try to write 10^4=(5*2)^4 and factor out that 2^4 ....

OpenStudy (lgbasallote):

this seems like one of them thingies with the 

OpenStudy (turingtest):

whatever, I'm terrible at these kinds of problems anyway :P

OpenStudy (anonymous):

This seems like getting out a TI-84 and abusing the parentheses button... Just type it all in....

OpenStudy (anonymous):

Oh, i just read your note about the calculator...

OpenStudy (kinggeorge):

My first instinct is to somehow use the fact that \(324=18^2\)

OpenStudy (anonymous):

I believe this is how you would do it by hand.

OpenStudy (anonymous):

two things, the nice factorization, and persistence. Had no idea it was going to work until that division near the end.

OpenStudy (blockcolder):

So in the end, it was just a telescoping product. Thanks a lot!